For the following equation, state the value of the discriminant and then describe the nature of the solutions.
16x^2-7x+4=0
What is the value of the discriminant___________?
Which one of the statements below is correct?
a. The equation has two imaginary solutions.
b. The equation has two real solutions.
c. The equation has one real solution?
Thanks for your help!

Answer by stanbon(75887)

For the following equation, state the value of the discriminant and then describe the nature of the solutions.
16x^2-7x+4=0
What is the value of the discriminant b^2-4ac
= 7^2-4*16*4 = -207
-----------------------------------
Which one of the statements below is correct?
a. The equation has two imaginary solutions because
the discriminant is negative.
Cheers,
Stan H.